OpenStudy (gabylovesyou):

8x2 = 24x is it in standard form ?

OpenStudy (gabylovesyou):

@nubeer

OpenStudy (gabylovesyou):

no its \[8x ^{2} = 24x\]

OpenStudy (gabylovesyou):

is the quadratic equation *** in standard form ?

OpenStudy (phi):

This shows you what standard form is http://www.mathsisfun.com/algebra/quadratic-equation.html you want an equation that = 0

OpenStudy (gabylovesyou):

so in standard form it will be 8x^2 + 24x = 0 ? or what

geerky42 (geerky42):

8x² - 24x = 0

geerky42 (geerky42):

Just subtract both sides by 24x.

OpenStudy (gabylovesyou):

ohh ok :P thnks

geerky42 (geerky42):

No problem.

OpenStudy (gabylovesyou):

how do i factor that equation ?

OpenStudy (gabylovesyou):

using (x )(x )

OpenStudy (anonymous):

look at both terms, determine what common factors they share, factor it out since its equal to zero, just solve

OpenStudy (phi):

you start with 8 x^2 = 24 x add -24x to both sides 8 x^2 -24 x = 24 x - 24 x simplify to get 8x^2 -24 x =0

OpenStudy (phi):

most people would immediately factor out an 8 from both terms

OpenStudy (phi):

in other words if you divide both sides by 8 you get x^2 - 3 x = 0/8 or simply x^2 - 3 x =0 you can now factor this. (factor out an x)

OpenStudy (gabylovesyou):

ok thank you